OpenStudy (anonymous):

Find g'(x) when g(x) = (2x^3 + 8x) ^ 1/2

OpenStudy (shamil98):

Are you familiar with the chain rule?

OpenStudy (shamil98):

Basically to derive this equation, you take the derivative of (2x^3 + 8x)^1/2 , with the power rule. and derive 2x^3 + 8x , seperately. and then multiply the two results

OpenStudy (shamil98):

The power rule. \[\large x^n = nx^{n-1}\]

OpenStudy (shamil98):

Do you understand thus far?

OpenStudy (anonymous):

no for the power rule how would I use that in this equation?

OpenStudy (shamil98):

(2x^3+8x)^1/2 becomes 1/2(2x^3+8x)^-1/2

OpenStudy (anonymous):

would I move the 1/2 infront?

OpenStudy (anonymous):

oh and then 1/2(2x^3 + 8x) ^-1/2 would multiple the 2x^3 +8x

OpenStudy (shamil98):

\[\large \frac{ (2x^3 + 8x)^{-1/2} }{ 2 } = \frac{ 1 }{ 2\sqrt{2x^3+8x} }\]

OpenStudy (shamil98):

Now, take the derivative of 2x^3 + 8 then multiply.

OpenStudy (anonymous):

6x^2+0 is the derivative?

OpenStudy (shamil98):

2x^3 + 8x** I made a typo

OpenStudy (anonymous):

6x^2+8

OpenStudy (shamil98):

Correct. Now you just multiply everything and simplify. \[\large g'(x) = \frac{ 6x^2 + 8 }{ 2\sqrt{2x^3+8x} } = \frac{ 3x^2+4 }{ \sqrt{2x^3+8x} }\]

OpenStudy (anonymous):

why is the \[2\sqrt{2x} thing on the bottom\]

OpenStudy (shamil98):

because \[\large x^{-n} = \frac{ 1 }{ x^n }\] \[\large \sqrt{x} = x^{1/2}\]

OpenStudy (anonymous):

oh thank you!

OpenStudy (shamil98):

Glad to be of assistance.